OpenStudy (anonymous):

Solve: log 4 (X+3) =2

OpenStudy (anonymous):

4 is the base??

OpenStudy (lalaly):

\[4^2=x+3\]\[x=13\]

OpenStudy (anonymous):

yes 4 is the base

OpenStudy (anonymous):

ok so do you follow that 4^log base 4 = 1?

OpenStudy (lalaly):

\[\log_b A=x\]becomes\[b^x=A\]